# Quantification of Interdependent Emotion Dynamics in Online Interactions

### v3

@@ Appendix E Robustness Checks
E.1   Comparison with Null Distribution
To assess whether the parameters estimated by our Hawkes model capture meaningful excitation dynamics rather than statistical noise, we conduct a null distribution analysis in the following. The null hypothesis posits that there is no self- or cross-excitation of emotions, meaning no emotional contagion or interaction occurs in the YouTube Live setting. Under this null, all entries of the $\boldsymbol{\alpha}$ and $\boldsymbol{\nu}$ matrices are set to zero. To simulate this scenario, we generated synthetic data from a six-dimensional Hawkes process in which the ground-truth $\boldsymbol{\alpha}$ and $\boldsymbol{\nu}$ matrices are all zero. In this setting, each emotion is independently driven by a constant baseline exogenous rate, $\mu_{0}^{e}=0.5$ for all $e$. We generated 100 synthetic datasets under this null specification, following the procedure described in Appendix D above. We then re-estimated the model parameters for each null dataset using the same log-likelihood maximization procedure as described in the Methods section. This yielded a distribution of parameter estimates that represents the null distribution, which effectively describes the expected behavior of the model in the absence of emotional excitation and video-driven influence.
@@ E.1   Comparison with Null Distribution
E.2   Robustness Check on Video Influence
Table E.1: Mean estimates of endogenous excitation parameters $\alpha_{e,f}$ and decay rates $\gamma^{e}$ (left), and exogenous influence parameters $\nu_{e,f}$ and spontaneous baselines $\mu_{0}^{e}$ (right) for different model specfications. The Baseline model corresponds to our main analysis, with results also reported in Figure 2. For all other models, we re-estimate the parameters using subsets of videos filtered by the specified keywords.
@@ E.2   Robustness Check on Video Influence
Table E.2: Statistics of emotion dynamics for different model specifications. The spectral radius $\rho$ of the branching ratio matrix $\alpha^{e,f}$ provides a global measure of self-excitation and cross-excitation. For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time, and obtain a fraction of endogenous versus exogenous intensity. For each emotion, we calculate and report the median endogenous vs. exogenous intensity ratio from the distribution of these ratios across videos. We obtain the median statistics of spontaneous vs. video-influenced intensity ratios for each emotion in the same manner. The Baseline model corresponds to our main analysis, with results also reported in Figure 3. For all other models, we report the results using subsets of videos filtered by the specified keywords. We notice that emotions are predominantly triggered endogenously.
@@
Table E.3: Mean estimates of endogenous excitation parameters $\alpha_{e,f}$ and decay rates $\gamma^{e}$ (left), and exogenous influence parameters $\nu_{e,f}$ and spontaneous baselines $\mu_{0}^{e}$ (right) for different model specfications. The Baseline model corresponds to our main analysis, with results also reported in Figure 2. Columns “Factor 0.5” and “Factor 2” show the estimated parameters when multiplying the maximum and median in the log-normal function by a factor of 0.5 and 2, respectively. Column “Power Law” shows the estimated parameters when parametrizing the video influence with a power law shape. The estimated parameters across different models are highly consistent with our main result.
@@
Table E.4: Statistics of emotion dynamics for different model specifications. The spectral radius $\rho$ of the branching ratio matrix $\alpha^{e,f}$ provides a global measure of self-excitation and cross-excitation. For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time, and obtain a fraction of endogenous versus exogenous intensity. For each emotion, we calculate and report the median endogenous vs. exogenous intensity ratio from the distribution of these ratios across videos. We obtain the median statistics of spontaneous vs. video-influenced intensity ratios for each emotion in the same manner. The Baseline model corresponds to our main analysis, with results also reported in Figure 3. For all other models, we report the results using subsets of videos filtered by the specified keywords. The relative influences are highly consistent with our main result across parametrizations of video influence. The endogenous influence is comparably larger than the exogenous influence, and the source of mother events is predominantly spontaneous user expressions.
@@
To examine the impact of video content on live chat discussions, we have parameterized and interpolated the prevalence of emotions in the video via sums of log-normal shapes (see Methods). Here, we test the robustness of our results with respect to changes in the log-normal parameters. We have defined the shape of the log-normal function by calibrating its mode (maximum value), and the median. We assumed that the intensity of each transcript peaks $2$ seconds after appearance such that the maximum of the log-normal function lies at $2$ seconds. We further assumed that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance such that the median is equal to $10$ seconds. Here we repeat our analysis while changing the log-normal function maximum and median by factors of 0.5 and 2 to test the robustness of our results. This effectively changes assumptions about the audience’s reaction time and memory of prior content. We report the results of these analyses in columns “Factor 0.5” and “Factor 2” of Tables E.3 and E.4. We observe that the results are qualitatively consistent with the results in the main text. This suggests that our results aren’t particularly sensitive to the parametrization of the video content.
@@
Human memory decays following an approximate power law [67]. The log-normal function resembles the shape of a power law over a limited range of the variable, which is all the larger, the larger is $\sigma$ [68]. We thus test our results with an alternative parametrization of the video influence in a power law shape where $s^{f}_{\tau_{j}}(t)=\frac{1}{(t-\tau_{j})^{c}}$. We report results in column “Power Law” of Tables E.3 and E.4 for $c=2.5$, and find that the results are qualitatively consistent with the results we obtain from the log-normal over a range of $c$ values from $2.5$ to $10$.
E.3   Robustness Check on Livechat Filtering
@@
Table E.5: Mean estimates of endogenous excitation parameters $\alpha_{e,f}$ and decay rates $\gamma^{e}$ (left), and exogenous influence parameters $\nu_{e,f}$ and spontaneous baselines $\mu_{0}^{e}$ (right) for different model specfications. The Baseline model corresponds to our main analysis, with results also reported in Figure 2. We conduct several robustness checks by varying key preprocessing steps and model configurations. We further conduct robustness checks using alternative data processing and modeling choices. In Model A, we include the full sample of live chat messages without filtering out messages sent before or after the video duration. In Model B, we exclude videos that are manually identified as centering around direct interactions between the live streamer and the audience. In the 4 Emotions specification, we restrict the analysis to a reduced set of emotions ($|\mathcal{E}|=4$), modeling only joy, anger, disgust, and sadness, while excluding fear and surprise. These alternative specifications help validate the robustness of our core findings.
@@
Table E.6: Statistics of emotion dynamics for different model specifications. The spectral radius $\rho$ of the branching ratio matrix $\alpha^{e,f}$ provides a global measure of self-excitation and cross-excitation. For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time, and obtain a fraction of endogenous versus exogenous intensity. For each emotion, we calculate and report the median endogenous vs. exogenous intensity ratio from the distribution of these ratios across videos. We obtain the median statistics of spontaneous vs. video-influenced intensity ratios for each emotion in the same manner. The Baseline model corresponds to our main analysis, with results also reported in Figure 3. In Model A, we include the full sample of live chat messages without filtering out messages sent before or after the video duration. In Model B, we exclude videos that are manually identified as centering around direct interactions between the live streamer and the audience. In the 4 Emotions specification, we restrict the analysis to a reduced set of emotions ($|\mathcal{E}|=4$), modeling only joy, anger, disgust, and sadness, while excluding fear and surprise.
@@
Nonetheless, we examine the robustness of our findings with respect to the filtering of live chat messages. We use an updated data sample without filtering out live chat messages that occurred before or after the official video runtime for each video. The resulting observation windows for each video span from the timestamp of the first live chat message to that of the last. The video influence at each time point is computed following the same procedure detailed in the Methods section. The results of this extended analysis are presented in column “Model A” of Tables E.5 and E.6. We find that including the full set of live chats does not alter our main conclusions, further reinforcing the robustness of our findings.
E.4   Robustness Check on Live Interactions
@@
To examine the sensitivity of our results to this assumption, we excluded all fully interactive videos (label = 2) and re-estimated the model on the remaining 241 videos. The results from this restricted sample are reported under “Model B” in Tables E.5 and E.6. We find that the estimates remain qualitatively consistent with our main findings.
Additionally, we note that if the audience does influence the live streamer, then the video content itself becomes partially endogenous to the system. This introduces a circular dynamic: chat influences video, which in turn influences chat. In such cases, the video—originally treated as an exogenous source—would, in fact, contain endogenous components. This feedback loop would further strengthen our core conclusion: endogenous dynamics, whether direct (chat-to-chat) or mediated via the video, play a dominant role in shaping emotional expression during live interactions.
E.5   Dynamics of 4 Emotions
Our main analysis models the commonly studied 6 basic emotions [50]. Here we test the robustness of our results by modeling $|\mathcal{E}|=4$ emotions: joy, anger, disgust, and sadness, removing fear and surprise because they have small excitatory patterns and low overall intensity. We obtain a sample of 288,581 live chat messages across 1,405 videos. Reducing the dimensionality of the multivariate Hawkes process reduces the number of parameters to be estimated from 84 to 40. We report our results for 4 emotions in column “4 Emotions” of Tables E.5 and E.6. We find that the dynamics of 4 emotions are qualitatively similar to our main results on 6 emotions (Figures 2 and 3 in the main text).
E.6   Dynamics of Live Chats with No Emotional Content
@@
Table E.7: The mean and standard deviation of estimated parameters for a univariate Hawkes Process of live chat messages with no emotion labels across 10 bootstrapped samples.
Figure E.1: The left plot shows the relative ratios of endogenous and exogenous influences for live chat messages with no emotion labels. The right plot shows the ratios of mother events from spontaneous expression and video influence for live chat messages with no emotion labels. We observe that the endogenous influence nonetheless dominates the dynamics of live chat messages with no emotion labels. However, the source of mother events is distributed more evenly between spontaneous expressions and video influence.
@@
The value of $S(t)$ is given by $S(t)=\sum_{\tau_{j}<t}s_{\tau_{j}}(t)$ where $\left\{\tau_{j}\right\}$ enumerates all times at which a subtitle first appears in the video. The values of $s_{\tau_{j}}(t)$ are calculated according to Equation (3) from the main text with $\mu=2.3$ and $\sigma=1.3$. The term $s_{\tau_{j}}(t)$ represents the time-varying intensity arising from a video subscript appearing in the video at time $\tau_{j}$. Similarly, we assume that $\phi$ is exponentially decaying where $\phi(t)=\alpha\frac{1}{\gamma}e^{-\frac{1}{\gamma}t}$ with decay time $\gamma$. Table E.7 shows the estimated parameters of Equation (10). We observe that live chat messages with no emotional content are nonetheless highly self-exciting but with a shorter direct memory than emotional live chats. Unlike our main result (Figure 3 in the main text), live chat messages with no emotional content are driven more equally by the video and user spontaneity, as shown in Figure E.1. This shows that while live discussions evolve largely endogenously, live chats with no emotional content are much more affected by videos. In contrast, emotions on online platforms are largely attributed to user expressions and interactions.

### v4

@@ Appendix 5 Robustness Checks
5.1  Comparison with Null Distribution
To assess whether the parameters estimated by our Hawkes model capture meaningful excitation dynamics rather than statistical noise, we conduct a null distribution analysis in the following. The null hypothesis posits that there is no self- or cross-excitation of emotions, meaning no emotional contagion or interaction occurs in the YouTube Live setting. Under this null, all entries of the $\boldsymbol{\alpha}$ and $\boldsymbol{\nu}$ matrices are set to zero. To simulate this scenario, we generated synthetic data from a six-dimensional Hawkes process in which the ground-truth $\boldsymbol{\alpha}$ and $\boldsymbol{\nu}$ matrices are all zero. In this setting, each emotion is independently driven by a constant baseline exogenous rate, $\mu_{0}^{e}=0.5$ for all $e$. We generated 100 synthetic datasets under this null specification, following the procedure described in Supplementary Note 4 above. We then re-estimated the model parameters for each null dataset using the same log-likelihood maximization procedure as described in the Methods section. This yielded a distribution of parameter estimates that represents the null distribution, which effectively describes the expected behavior of the model in the absence of emotional excitation and video-driven influence.
@@ 5.1  Comparison with Null Distribution
5.2  Robustness Check on Video Influence
Supplementary Table 2: Mean estimates of endogenous excitation parameters $\alpha_{e,f}$ and decay rates $\gamma^{e}$ (left), and exogenous influence parameters $\nu_{e,f}$ and spontaneous baselines $\mu_{0}^{e}$ (right) for different model specfications. The Baseline model corresponds to our main analysis, with results also reported in Figure 2 in the main text. For all other models, we re-estimate the parameters using subsets of videos filtered by the specified keywords.
@@ 5.2  Robustness Check on Video Influence
Supplementary Table 3: Statistics of emotion dynamics for different model specifications. The spectral radius $\rho$ of the branching ratio matrix $\alpha^{e,f}$ provides a global measure of self-excitation and cross-excitation. For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time, and obtain a fraction of endogenous versus exogenous intensity. For each emotion, we calculate and report the median endogenous vs. exogenous intensity ratio from the distribution of these ratios across videos. We obtain the median statistics of spontaneous vs. video-influenced intensity ratios for each emotion in the same manner. The Baseline model corresponds to our main analysis, with results also reported in Figure 3 in the main text. For all other models, we report the results using subsets of videos filtered by the specified keywords. We notice that emotions are predominantly triggered endogenously.
@@
Supplementary Table 4: Mean estimates of endogenous excitation parameters $\alpha_{e,f}$ and decay rates $\gamma^{e}$ (left), and exogenous influence parameters $\nu_{e,f}$ and spontaneous baselines $\mu_{0}^{e}$ (right) for different model specfications. The Baseline model corresponds to our main analysis, with results also reported in Figure 2 in the main text. Columns Factor 0.5 and Factor 2 show the estimated parameters when multiplying the maximum and median in the log-normal function by a factor of 0.5 and 2, respectively. Column Power Law shows the estimated parameters when parametrizing the video influence with a power law shape. The estimated parameters across different models are highly consistent with our main result.
@@
Supplementary Table 5: Statistics of emotion dynamics for different model specifications. The spectral radius $\rho$ of the branching ratio matrix $\alpha^{e,f}$ provides a global measure of self-excitation and cross-excitation. For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time, and obtain a fraction of endogenous versus exogenous intensity. For each emotion, we calculate and report the median endogenous vs. exogenous intensity ratio from the distribution of these ratios across videos. We obtain the median statistics of spontaneous vs. video-influenced intensity ratios for each emotion in the same manner. The Baseline model corresponds to our main analysis, with results also reported in Figure 3 in the main text. For all other models, we report the results using subsets of videos filtered by the specified keywords. The relative influences are highly consistent with our main result across parametrizations of video influence. The endogenous influence is comparably larger than the exogenous influence, and the source of mother events is predominantly spontaneous user expressions.
@@
To examine the impact of video content on live chat discussions, we have parameterized and interpolated the prevalence of emotions in the video via sums of log-normal shapes (see Methods). Here, we test the robustness of our results with respect to changes in the log-normal parameters. We have defined the shape of the log-normal function by calibrating its mode (maximum value), and the median. We assumed that the intensity of each transcript peaks $2$ seconds after appearance such that the maximum of the log-normal function lies at $2$ seconds. We further assumed that 50% of the emotion intensity for each transcript is manifested within $10$ seconds of transcript appearance such that the median is equal to $10$ seconds. Here we repeat our analysis while changing the log-normal function maximum and median by factors of 0.5 and 2 to test the robustness of our results. This effectively changes assumptions about the audience’s reaction time and memory of prior content. We report the results of these analyses in columns Factor 0.5 and Factor 2 of Supplementary Tables 4 and 5. We observe that the results are qualitatively consistent with the results in the main text. This suggests that our results aren’t particularly sensitive to the parametrization of the video content.
@@
Human memory decays following an approximate power law [64]. The log-normal function resembles the shape of a power law over a limited range of the variable, which is all the larger, the larger is $\sigma$ [65]. We thus test our results with an alternative parametrization of the video influence in a power law shape where $s^{f}_{\tau_{j}}(t)=\frac{1}{(t-\tau_{j})^{c}}$. We report results in column Power Law of Supplementary Tables 4 and 5 for $c=2.5$, and find that the results are qualitatively consistent with the results we obtain from the log-normal over a range of $c$ values from $2.5$ to $10$.
5.3  Robustness Check on Livechat Filtering
@@
Supplementary Table 6: Mean estimates of endogenous excitation parameters $\alpha_{e,f}$ and decay rates $\gamma^{e}$ (left), and exogenous influence parameters $\nu_{e,f}$ and spontaneous baselines $\mu_{0}^{e}$ (right) for different model specfications. The Baseline model corresponds to our main analysis, with results also reported in Figure 2 in the main text. We conduct several robustness checks by varying key preprocessing steps and model configurations. We further conduct robustness checks using alternative data processing and modeling choices. In Model A, we include the full sample of live chat messages without filtering out messages sent before or after the video duration. In Model B, we exclude videos that are manually identified as centering around direct interactions between the live streamer and the audience. In the 4 Emotions specification, we restrict the analysis to a reduced set of emotions ($|\mathcal{E}|=4$), modeling only joy, anger, disgust, and sadness, while excluding fear and surprise. These alternative specifications help validate the robustness of our core findings.
@@
Supplementary Table 7: Statistics of emotion dynamics for different model specifications. The spectral radius $\rho$ of the branching ratio matrix $\alpha^{e,f}$ provides a global measure of self-excitation and cross-excitation. For each video, we calculate the average ratio of endogenous (exogenous) intensity to the total intensity across time, and obtain a fraction of endogenous versus exogenous intensity. For each emotion, we calculate and report the median endogenous vs. exogenous intensity ratio from the distribution of these ratios across videos. We obtain the median statistics of spontaneous vs. video-influenced intensity ratios for each emotion in the same manner. The Baseline model corresponds to our main analysis, with results also reported in Figure 3 in the main text. In Model A, we include the full sample of live chat messages without filtering out messages sent before or after the video duration. In Model B, we exclude videos that are manually identified as centering around direct interactions between the live streamer and the audience. In the 4 Emotions specification, we restrict the analysis to a reduced set of emotions ($|\mathcal{E}|=4$), modeling only joy, anger, disgust, and sadness, while excluding fear and surprise.
@@
Nonetheless, we examine the robustness of our findings with respect to the filtering of live chat messages. We use an updated data sample without filtering out live chat messages that occurred before or after the official video runtime for each video. The resulting observation windows for each video span from the timestamp of the first live chat message to that of the last. The video influence at each time point is computed following the same procedure detailed in the Methods section. The results of this extended analysis are presented in column Model A of Supplementary Tables 6 and 7. We find that including the full set of live chats does not alter our main conclusions, further reinforcing the robustness of our findings.
5.4  Robustness Check on Live Interactions
@@
To examine the sensitivity of our results to this assumption, we excluded all fully interactive videos (label = 2) and re-estimated the model on the remaining 241 videos. The results from this restricted sample are reported under “Model B” in Supplementary Tables 6 and 7. We find that the estimates remain qualitatively consistent with our main findings. Moreover, the numerical differences between the baseline model and Model B generally fall well within one standard deviation of our bootstrapped parameter estimates, suggesting that the observed variation is likely due to statistical noise. To formally verify this, we conducted two-sample t-tests comparing the parameter distributions of the baseline and restricted models across all six emotions and for all parameters, $\mu^{e}_{0}$, $\nu^{e,f}$, $\alpha^{e,f}$, and $\gamma^{e}$. For all cases, the resulting $p$-values were greater than $0.1$, indicating that we cannot reject the null hypothesis that the two sets of parameter estimates are drawn from the same underlying distribution. These results confirm that the presence or absence of fully interactive videos does not meaningfully affect our overall conclusions.
Additionally, we note that if the audience does influence the live streamer, then the video content itself becomes partially endogenous to the system. This introduces a circular dynamic: chat influences video, which in turn influences chat. In such cases, the video - originally treated as an exogenous source - would, in fact, contain endogenous components. This feedback loop would further strengthen our core conclusion: endogenous dynamics, whether direct (chat-to-chat) or mediated via the video, play a dominant role in shaping emotional expression during live interactions.
5.5  Dynamics of 4 Emotions
Our main analysis models the commonly studied 6 basic emotions [47]. Here we test the robustness of our results by modeling $|\mathcal{E}|=4$ emotions: joy, anger, disgust, and sadness, removing fear and surprise because they have small excitatory patterns and low overall intensity. We obtain a sample of 288,581 live chat messages across 1,405 videos. Reducing the dimensionality of the multivariate Hawkes process reduces the number of parameters to be estimated from 84 to 40. We report our results for 4 emotions in column 4 Emotions of Supplementary Tables 6 and 7. We find that the dynamics of 4 emotions are qualitatively similar to our main results on 6 emotions (Figures 2 and 3 in the main text).
5.6  Dynamics of Live Chats with No Emotional Content
@@
Supplementary Table 8: The mean and standard deviation of estimated parameters for a univariate Hawkes Process of live chat messages with no emotion labels across 10 bootstrapped samples.
Supplementary Figure 7: The left plot shows the relative ratios of endogenous and exogenous influences for live chat messages with no emotion labels. The right plot shows the ratios of mother events from spontaneous expression and video influence for live chat messages with no emotion labels. We observe that the endogenous influence nonetheless dominates the dynamics of live chat messages with no emotion labels. However, the source of mother events is distributed more evenly between spontaneous expressions and video influence.
@@
The value of $S(t)$ is given by $S(t)=\sum_{\tau_{j}<t}s_{\tau_{j}}(t)$ where $\left\{\tau_{j}\right\}$ enumerates all times at which a subtitle first appears in the video. The values of $s_{\tau_{j}}(t)$ are calculated according to Equation (3) from the main text with $\mu=2.3$ and $\sigma=1.3$. The term $s_{\tau_{j}}(t)$ represents the time-varying intensity arising from a video subscript appearing in the video at time $\tau_{j}$. Similarly, we assume that $\phi$ is exponentially decaying where $\phi(t)=\alpha\frac{1}{\gamma}e^{-\frac{1}{\gamma}t}$ with decay time $\gamma$. Supplementary Table 8 shows the estimated parameters of Equation (15). We observe that live chat messages with no emotional content are nonetheless highly self-exciting but with a shorter direct memory than emotional live chats. Unlike our main result (Figure 3 in the main text), live chat messages with no emotional content are driven more equally by the video and user spontaneity, as shown in Supplementary Figure 7. This shows that while live discussions evolve largely endogenously, live chats with no emotional content are much more affected by videos. In contrast, emotions on online platforms are largely attributed to user expressions and interactions.